# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 7.2. End-to-end performance comparisons
Observation 1. SPECIAL significantly outperforms all baseline systems in query latency, achieving performance gains of up to 3618.3×3618.3$\times$3618.3 × for linear queries, 114×114$\times$114 × for binary joins, and 80.3×80.3$\times$80.3 × for multi-joins. Figure 5 shows significant performance gains of SPECIAL in contrast to Shrinkwrap and SMCQL. For linear queries, SPECIAL exhibits substantial performance gains, achieving up to 3618.3×3618.3$\times$3618.3 × improvement (Q2), primarily due to its index-based fast data access. This approach not only reduces large I/O costs but also eliminates the need for secure computation by enabling direct data retrieval via DP indexes. We observe that in binary joins, SPECIAL has a less pronounced advantage over Shrinkwrap. This is because Shrinkwrap do not have sub-optimal planning issue when dealing with binary join—as there is only one join order. Hence, the performance gains for SPECIAL in Q3 and Q4 stem solely from fast data access and the efficient (MX)Join method. However, for more complex multi-way joins, SPECIAL’s advantage becomes more pronounced again due to its ability to select the optimal join order. For instance, Shrinkwrap shows more than 80×80$\times$80 × slow down in query latency than SPECIAL in group Q7.
@@
Observation 2. SPECIAL shows profound improvement in memory usage (Figure 6) against baseline systems, especially in complex multi-way joins. This is primarily due to two factors: First, the (MX)JOIN used by SPECIAL is more memory-efficient compared to the joins implemented by Shrinkwrap and SMCQL. Second, SPECIAL’s capability to identify optimal execution plans significantly reduces total intermediate sizes, which is particularly beneficial for complex joins that suffer from sub-optimal or exhaustive padding in other systems. To better understand the substantial improvements SPECIAL achieves—for instance, up to 928.2×928.2$\times$928.2 × over Shrinkwrap and more than $10^{5}\times$ over SMCQL—we will zoom into a specific query, Q6, and compare the detailed execution plans of the three systems. The choice of Q6 is strategic because its complexity sufficiently highlights the differences in execution plans, yet it remains simple enough for clear visual representation. Our comparison features four execution plans: a plaintext optimal plan derived directly from the logical plan using actual cardinalities, illustrating the ground truth optimal execution; a hypothetical SMCQL plan (since Q6 cannot be completed by SMCQL so we project the cardinalities); and two actual execution plans from our experiments with Shrinkwrap and SPECIAL. The detailed comparisons and observations are summarized in Figure 4.
7.3. Privacy comparisons
@@
Continue to address Question-1, we now compare the privacy guarantees between Shrinkwrap and SPECIAL, with the results detailed in Figure 7. As privacy loss is independent of query types, our analysis will focus on comparing the privacy composition curves under continual query answering for both Shrinkwrap and SPECIAL. We consider two privacy composition models: advanced composition (Adv.)(Dwork et al., 2014) and concentrated DP composition (CDP)(Bun and Steinke, 2016).
@@
Observation 3. Under continual query answering, SPECIAL demonstrates significantly lower privacy loss compared to Shrinkwrap, achieving up to 89.01×89.01$\times$89.01 × and 38.91×38.91$\times$38.91 × improvements in the Adv. and CDP modes, respectively. The privacy loss of SPECIAL is bounded to the initial synopsis release stage, so continual query answering does not incur additional privacy loss. In contrast, Shrinkwrap’s privacy loss accumulates over time as each new query allocates a fresh privacy budget. Consequently, its privacy loss exhibits a logarithmic growth, as shown in Figure 7. This accumulation can result in significant privacy degradation when processing a large number of queries. For example, answering 100 queries in Shrinkwrap could result in a privacy loss of $\epsilon>100$ in Adv. and $\epsilon\approx 60$ in CDP., respecitvely, even if each query only uses a small privacy budget of $\epsilon=1.5$. As such, SPECIAL demonstrate significant improve in privacy guarantees towards SOTA DPSCA.
@@ 7.4. Privacy efficiency tradeoffs
We address Question-2 by evaluating SPECIAL at various privacy levels and observing the performance impacts. Specifically, we maintain $\delta$ constant while varying $\epsilon$ from $0.1$ to $10$ and assess the performance across default testing queries (Q2, Q4, Q8). The results are shown in Figure 8 and Figure 9.
Figure 8. Privacy vs. Performance (query time)
Figure 9. Privacy vs. Performance (memory)
@@
Observation 5. Different queries exhibit varying performance sensitivities to changes in privacy levels. Multi-join queries are significantly affected when privacy levels change. For instance, the performance overhead of Q8 at $\epsilon=0.1$ is more than 3×3$\times$3 × higher than at $\epsilon=1$. In contrast, linear query (Q2) and binary join (Q4) overhead increases to only about 1.5×1.5$\times$1.5 × under the same privacy settings. This difference stems from the error amplification inherent in multi-join operations, which frequently reuse synopses and require multiple cardinality estimations. Therefore, practitioners dealing with complex multi-join workloads may need careful fine-tune privacy parameters, as performance impacts can be substantial.
@@ 7.5. Scaling experiments
To address Question-3, we stress SPECIAL with scaled workloads by duplicating the raw financial dataset to sizes of 2×2$\times$2 ×, 4×4$\times$4 ×, and 8×8$\times$8 ×. We then evaluate the default testing queries on these scaled datasets to assess SPECIAL’s efficiency. The results are shown in Figure 10.
Figure 10. End-to-end comparison privacy loss.
Observation 6. SPECIAL demonstrates significant potential to scale up to multi-million records, even when handling complex 5-way joins. Figure 10 shows SPECIAL’s effective scaling: up to 8×8$\times$8 × data for linear queries and binary joins, and up to 4×4$\times$4 × data for complex 5-way joins like Q8. For instance, Q2 can be completed within 290ms under 8×8$\times$8 ×, and in fact, since selection is bypassed (due to index access), thus the cost is mainly on I/O costs. Q4 finishes in 289 minutes at the same scale 8×8$\times$8 ×, while the more complex 5-way join Q8 takes less than 280 minutes for 4×4$\times$4 × data. As a reference, Shrinkwrap would require over 1035 minutes to complete Q8 even with unscaled data (Figure 5). This result can, from another angle, showcase the significant performance improvements of SPECIAL over current SOTA DPSCA designs.
8. Discussion
In this section, we discuss potential extensions to accommodate broader data models and possible adaptations of SPECIAL designs to other secure primitives.
Supporting growing data. In general SPECIAL considers read-intensive analytical workloads and and a static data model, which is the same as current SOTA SCA efforts (Poddar et al., 2021; Bater et al., 2018; Liagouris et al., 2023; Volgushev et al., 2019). In this section, we will briefly discuss how SPECIAL could be extended to support the growing data model where new records are continuously added to the logical database (Wang et al., 2022, 2021; Zhang et al., 2023). One possible strategy involves segmenting the large logical database into discrete, non-overlapping time intervals, each represented by its own database with independent synopses. Specifically, for new incoming data, SPECIAL will initially cache it until a designated time interval concludes. After this period, SPECIAL will batch all the cached data into a new database corresponding to the past time interval and release synopses for this database. As the data within these intervals is disjoint, queries can be easily partitioned and independently optimized and executed across corresponding databases in parallel, before being aggregated. Note that prior to releasing new synopses, all cached data will be processed using standard SCA processing (e.g. SMCQL). Since each synopses release covers only newly inserted data, the total privacy loss adheres to parallel composition (Dwork et al., 2014), and thus the bounded privacy loss still hold.
Adaptability to other secure primitives. Beyond MPC-based SCA, the key design insights from SPECIAL are transferable to other secure outsourced databases such as the TEE-based analytical systems (Qiu et al., [n.d.]; Qin et al., 2022; Eskandarian and Zaharia, 2017; Zheng et al., 2017) or searchable encryption based private key-value store (Amjad et al., 2021; Groce et al., 2019; Chen et al., 2018). This is because SPECIAL’s core components, including synopses management, indexing, query planning, and compaction estimation, can all be executed without relying on secure computation primitives. For example, in TEE databases, one may conduct a similar query planning outside the TEE, with the optimal plan then executed within the TEE for fast oblivious processing. Similarly, SPECIAL’s design principles could be applied to searchable encryption to construct private indexes in advance on sorted data, enable fast and private information retrieval.
@@ 9. Related Work
SCA systems. Two main approaches exist for designing MPC-based SCA systems. The first is peer-to-peer paradigm (Aggarwal et al., 2005; Mouris et al., 2024; Poddar et al., 2021; Volgushev et al., 2019; Wang and Yi, 2021), where the goal is to improve efficiency by decomposing analytical queries and pushing them to data owners, so that they can either directly process in clear or running MPCs across a smaller group of parties. Unfortunately, this approach burdens data owners with heavy computational overheads, particularly for complex operations like joins. Moreover, given real-world data owners often lack robust computing resources and reliable service capabilities, these methods are difficult to scale and promise consistent SCA services to external analysts. On the other hand, the server-aided-MPC model (Wang et al., 2021; Bater et al., 2018, 2020; Wang et al., 2022; Zhang et al., 2023; Roy Chowdhury et al., 2020; Kamara et al., 2011; Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023; Mohassel and Zhang, 2017) offers a more practical solution. It allows data owners to outsource both expensive MPC computations and secure data storage to a set of capable cloud service providers (CSPs), who then jointly evaluate MPC to provide reliable SCA services. For the SPECIAL prototype, we focus on the server-aided-MPC model under a strong corruption assumption, allowing up to $n-1$ of $n$ servers to be compromised and collate with others. While, as we discussed before, SPECIAL’s core design principles are protocol-agnostic. This enables interoperability with various MPC models, including the peer-to-peer settings or a weaker corruption setting (for performance purpose) where a supermajority of servers need to be honest (Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023).
@@
DP leakages. Leakage-abuse attacks (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016), exploit data-dependent processing patterns, are persistent threats to SCA systems. To mitigate these risks, oblivious computation techniques (Keller and Scholl, 2014; Pinkas and Reinman, 2010; Stefanov et al., 2011; Chang et al., 2022; Mishra et al., 2018; Crooks et al., 2018; Jönsson et al., 2011; Sasy et al., 2022; Asharov et al., 2020; Wang et al., 2014; Pettai and Laud, 2015; Wang et al., 2023; Tinoco et al., 2023) have become the de facto solution—processing data such that the entire execution transcript (e.g. memory traces and read/write volumes) is padded to a worst-case constant. While these techniques ensure the strongest privacy guarantees by eliminating data-dependent leakages, they also introduce a fundamental contention with modern database optimizations, which often rely heavily on data-dependent operations (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022). To this end, many recent efforts seek a practical balance in the privacy-performance trade-off by allowing controlled leakage under DP (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Groce et al., 2019; He et al., 2017; Zhang et al., 2023; Qin et al., 2022; Chan et al., 2022; Wagh et al., 2016; Chu et al., 2021; Qiu et al., [n.d.]; Patel et al., 2019). However, a common issue of these approaches is unbounded privacy loss. While some works propose to address this (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., [n.d.]), their approaches are restricted to only simple linear queries. Furthermore, existing DP leakage designs also lack lossless guarantees for complex query processing. The inherent randomness of DP noises (Wang et al., 2021; Groce et al., 2019; Wang et al., 2022) and uncertainty in complex join sensitivities (Chu et al., 2021; Bater et al., 2018; Wang et al., 2022) often lead to improper output compaction and lose tuples. SPECIAL addresses all these limitations together, and to our knowledge, is the first system to ensure both bounded privacy and lossless results for securely processing complex queries.
@@
SCA query planning. Query planning (Silberschatz et al., 2011) is crucial in conventional databases. A few studies (Bater et al., 2017; Volgushev et al., 2019; Poddar et al., 2021; Liagouris et al., 2023) explored query planning within SCA frameworks. Unlike conventional planners, which exploit size disparities across different execution plans to choose the most efficient one with minimized plan sizes (Elmasri et al., 2020; Harmouch and Naumann, 2017; Han et al., 2021), these methods typically prohibit the use of such data-dependent information. Consequently, they primarily rely on data-independent metrics for planning. For example, (Bater et al., 2017; Liagouris et al., 2023; Volgushev et al., 2019) optimize based on the MPC execution costs, while (Poddar et al., 2021) focuses on decomposing large MPC tasks into smaller 2/3PC ones. Since these methods do not optimize or compact intermediate sizes, significant query execution overhead from these large sizes remains evident, especially when processing complex multi-joins (Bater et al., 2018; Liagouris et al., 2023). Shrinkwrap (Bater et al., 2018) uses DP leakage models and allows planner to optimize query intermediate sizes in a privacy-preserving manner. However, it cannot estimate cardinalities before runtime, mandating a random selection of plan structures (e.g. join order) before execution. Consequently, planning focuses solely on optimally allocating privacy budgets across the chosen plan, which often result in suboptimal query executions. In contrast, SPECIAL offers an advanced query planner capable of pre-estimating plan sizes and comparing costs among different plan structures before execution.

### v2

@@ 7.2. End-to-end comparisons
Observation 1. SPECIAL outperforms Shrinkwrap and SMCQL in query latency across all benchmarks, reaching up to $3618.3\times$ for linear queries, $114\times$ for binary joins, and $80.3\times$ for multi-joins. Figure 4 shows the comparison results in query time. First, SPECIAL shows significant speedups for linear queries, reaching up to $3618.3\times$ (FQ2). This large performance gain is mainly attributed to its index-based fast data access. By directly fetching private data through DP indexes, SPECIAL eliminate substantial I/O costs (e.g., sequential reads) and bypass the need for secure computations (e.g., oblivious filter). Second, we observe that in binary joins, SPECIAL has a less pronounced advantage over Shrinkwrap. This is because binary joins have a single join order, eliminating the potential for join ordering problems (where different join orders lead to significantly different performance). Consequently, even though Shrinkwrap does not optimize join orders, it doesn’t experience efficiency losses in this scenario. However, for more complex multi-way joins, SPECIAL’s advantage becomes more pronounced again. For instance, more than $80\times$ speedup in FQ7. This is because SPECIAL can pre-select efficient join orders before runtime.
@@
Observation 2. SPECIAL shows profound improvement in memory usage (Figure 5) against baseline systems, especially in complex multi-way joins. This is primarily due to two factors: First, the (MX)JOIN used by SPECIAL is more memory-efficient compared to the joins implemented by Shrinkwrap and SMCQL. Second, SPECIAL’s capability to identify optimal execution plans significantly reduces total intermediate sizes, which is particularly beneficial for complex joins that suffer from sub-optimal or exhaustive padding in other systems. To better understand the substantial improvements SPECIAL achieves—for instance, up to $928.2\times$ over Shrinkwrap and more than $10^{5}\times$ over SMCQL—we will zoom into a specific query, FQ6, and compare the detailed execution plans of the three systems. The choice of FQ6 is strategic because its complexity sufficiently highlights the differences in execution plans, yet it remains simple enough for clear visual representation. Our comparison features four execution plans: a plaintext optimal plan, illustrating the ground truth optimal execution; a hypothetical SMCQL plan (with projected cardinalities); and two actual execution plans from our experiments with Shrinkwrap and SPECIAL. The detailed comparisons and observations are summarized in Figure 3.
@@
To continue address Question-1, we now compare the privacy guarantees of SPECIAL with the baseline systems (Figure 6). Specifically, we focus on comparing cumulative privacy loss in multiple query answering, w.r.t. two composition models: advanced composition (Adv.)(Dwork et al., 2014) and concentrated composition (CDP.)(Bun and Steinke, 2016).
@@
Observation 3. Under continual query answering, SPECIAL demonstrates significantly lower privacy loss compared to Shrinkwrap, achieving up to $89.01\times$ and $38.91\times$ improvements in the Adv. and CDP modes, respectively. The privacy loss of SPECIAL is bounded to the initial synopsis release stage, so continual query answering does not incur additional privacy loss. In contrast, Shrinkwrap’s privacy loss accumulates over time as each new query allocates a fresh privacy budget. Consequently, its privacy loss exhibits a logarithmic growth, as shown in Figure 6. This accumulation can result in significant privacy degradation when processing a large number of queries. For example, answering 100 queries in Shrinkwrap could result in a privacy loss of $\epsilon>100$ in Adv. and $\epsilon\approx 60$ in CDP., respecitvely, even if each query only uses a small privacy budget of $\epsilon=1.5$. As such, SPECIAL demonstrate significant improve in privacy guarantees towards SOTA DPSCA. Even when compared to standard SCA (e.g., SMCQL) with no privacy loss due to exhaustive padding, our system incurs only a small and fixed privacy cost (e.g. $\epsilon=1.5$ per table) while delivering substantial performance gains.
@@ 7.3. Privacy efficiency tradeoffs
We address Question-2 by evaluating SPECIAL at various privacy levels. Specifically, we maintain $\delta$ constant while varying $\epsilon$ from $0.1$ to $10$ and assess the performance across default testing queries. The results are shown in Figure 7.
Figure 7. Privacy vs. Performance trade-offs
@@
7.4. Synopses impacts micro benchmarks
We continue to address Question-2 to explore how synopses scenarios may affect SPECIAL’s performance. Specifically, we study two key settings: (i) How different bin numbers (BinNum) in synopses can impact the efficiency of (IDX)JOIN, and (ii) how synopsis coverage levels for a single query can affect its overall execution. We will conduct micro-benchmarks for a thorough investigation. Note that simulate different synopses scenario on both HealthLNK and Financial benchmarks can be challenging (e.g., joins typically occur on the same key, so it is hard to simulate partial coverage), hence, to better control experimental variables and accurately assess impacts, we will now use synthetic data and workloads.
Figure 8. BinNum experiments.
We first study how the BinNums impact the performance of (IDX)JOIN. To study this, we synthesize two join queries on fixed input data, generate join key synopses with varying bin numbers (2 to 64), and measure the performance of (IDX)JOIN in processing the queries. The results are shown in Figure 8.
Observation 6. The running time of (IDX)JOIN initially decreases but then increases as the BinNum grows. The memory usage consistently increases. (IDX)JOIN partitions larger joins into smaller sub-joins, but since we rely on DP indexes for partitioning, each sub-join inevitably includes additional dummy data. This increased plan size directly translates to higher memory usage and will grow when BinNum increases (more noises in DP indexes). On the other hand, partitioning large joins into smaller ones can enhance join efficiency, which is why we initially observe a decrease in execution time as BinNums increase (e.g., 2 to 8). However, the trade-off arises when the number of bins becomes excessive (e.g., >8>8). The overhead of handling the increased dummy data starts to outweigh the benefits gained from partitioning. At this point, the performance improvement plateaus and starts to degrade as the system struggles with the inflated plan sizes.
Table 3. Synopses coverage experiments.
Synopses Coverage Time (ms) Improv. Mem. (bytes) Improv.
No coveragee 1831703 baseline 129048576 baseline
Join Key
[…]
* We synthesize a random selectivity between (0, 0.33) for each filter operation.
We now study how synopses coverage impacts query performance. We synthesize and test a 4-way join query under two controlled scenarios: (i) JK coverage: We focus on varying the level of JK coverage, starting from 1 out of 3 JKs to full coverage, while ensuring no filter synopses are present. We then measure how this impacts query performance; and (ii) Filter coverage: We maintain full JK coverage and change the coverage of filter synopses on the query’s input tables, ranging from 1 out of 4 inputs to full coverage. This allows us to examine the isolated effect of filter synopsis coverage on performance. Results are in Table 3.
[…]
@@ 7.5. Scaling experiments
To address Question-3, we stress SPECIAL with two types of scaled workloads: (i) scaled data: we duplicate the raw dataset to sizes of $2\times$, $4\times$, and $8\times$ and evaluate default testing queries; (ii) scaled query complexity: We use standard inputs, but simulate complex multi joins (up to 9-way) by chaining together multiple join workloads. The results are shown in Figure 9 and Table 4.
Figure 9. Scaling data experiments.
Observation 8. SPECIAL shows large potential to scale up to multi-million data, even for complex 5-way joins. Figure 9 shows SPECIAL’s effective scaling: up to $8\times$ data for linear queries and binary joins, and up to $4\times$ data for complex 5-way joins like Q8. For instance, Q2 can be completed within 290ms under $8\times$, and in fact, since selection is bypassed (due to index access), thus the cost is mainly on I/O costs. Q4 finishes in 289 minutes at the same scale $8\times$, while the more complex 5-way join Q8 takes less than 280 minutes for $4\times$ data. As a reference, Shrinkwrap would require over 1035 minutes to complete Q8 even with unscaled data.
[…]
Table 4. Scaling query complexity experiments
Join scale Query time (ms) Improv. Memory (bytes) Improv.
Shrinkwrap Q6 58342782 baseline 3354165360 baseline
[…]
@@ 8. Related Work
SCA systems. Two main approaches exist for designing MPC-based SCA systems. The first is peer-to-peer (P2P) paradigm (Aggarwal et al., 2005; Mouris et al., 2024; Poddar et al., 2021; Volgushev et al., 2019; Wang and Yi, 2021), where the goal is to improve efficiency by decomposing analytical queries and pushing them to data owners, so that they can either directly process in clear or running MPCs across a small group of parties. Unfortunately, this approach imposes large overhead on data owners, especially for complex operations like joins. Given that real-world data owners often lack robust computing resources and service capabilities, the P2P paradigm is hard to scale and support reliable SCA services to external analysts. The other paradigm is the server-aided-MPC model (Wang et al., 2021; Bater et al., 2018, 2020; Wang et al., 2022; Zhang et al., 2023; Roy Chowdhury et al., 2020; Kamara et al., 2011; Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023; Mohassel and Zhang, 2017). This model allows data owners to outsource both expensive MPC computations and secure data storage to a set of capable servers, which can then jointly evaluate MPC to provide reliable SCA services. SPECIAL is built upon the server-aided-MPC model and under a strong “all but one”corruption assumption. Moreover, SPECIAL’s core design is protocol-agnostic, which allows interoperability with various MPC models, including the P2P or a weaker corruption where a supermajority of servers need to be honest (Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023).
@@
DP leakages. Leakage-abuse attacks (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016), exploit data-dependent processing patterns, are persistent threats to SCA systems. To mitigate these risks, oblivious computation (Keller and Scholl, 2014; Pinkas and Reinman, 2010; Stefanov et al., 2011; Chang et al., 2022; Mishra et al., 2018; Crooks et al., 2018; Jönsson et al., 2011; Sasy et al., 2022; Asharov et al., 2020; Wang et al., 2014; Pettai and Laud, 2015; Wang et al., 2023; Tinoco et al., 2023) have become the de facto solution. While this technique ensures the strongest privacy guarantees by eliminating any data-dependent leakages, it also introduces a fundamental contention with modern database optimizations, which often rely heavily on data-dependent operations (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022). To this end, many recent efforts seek a practical balance in the privacy-performance trade-off by allowing controlled leakage under DP (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Groce et al., 2019; He et al., 2017; Zhang et al., 2023; Qin et al., 2022; Chan et al., 2022; Wagh et al., 2016; Chu et al., 2021; Qiu et al., 2023; Patel et al., 2019). However, a common issue of these approaches is unbounded privacy loss. While some works propose to address this (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., 2023), their approaches are restricted to only simple linear queries. SPECIAL addresses all these limitations together, and to our knowledge, is the first SCA system that can simultaneously ensure both bounded privacy and lossless results for complex SPJA queries.
@@
SCA query planning. Query planning (Silberschatz et al., 2011) is crucial in conventional databases. Conventional planners can exploit size disparities across different query plans to choose efficient ones with smaller sizes (Elmasri et al., 2020; Harmouch and Naumann, 2017; Han et al., 2021). However, such techniques use data-dependent information and are typically prohibited in SCA. A handful of studies (Bater et al., 2017; Volgushev et al., 2019; Poddar et al., 2021; Liagouris et al., 2023) that explore query planning within SCA frameworks primarily rely on data-independent metrics for planning, which usually lead to only moderate optimizations. Shrinkwrap (Bater et al., 2018) introduced a private planning method that optimally compacts intermediate sizes by efficiently allocating privacy budgets to minimize dummy data. However, it cannot pre-determine an optimal join order. SPECIAL offers an advanced query planner capable of pre-estimating intermediate sizes and comparing execution costs among different plan structures before runtime.